# Homogenization with Guaranteed Bounds via Primal-Dual Physically Informed Neural Networks

### v1

@@ 3 Physically Informed Neural Networks for Homogenization
To show how existing PINN homogenization approaches can be improved by introducing the dual form training loss, we adopt the PINN architecture introduced by Jiang et al. (2023) (Fig. 1), which was demonstrated to produce a good approximation of the solution of (1) for a wide range of different material functions with smooth coefficients, see Fig. 1. This network is characterized by a unique periodic layer that employs learnable cosine functions. This design choice ensures that the network’s output is inherently periodic, thereby eliminating the need to incorporate a boundary condition term into the training loss.
The network takes as input the microscopic coordinates $x_{1}$ and $x_{2}$ and produces output in one of two forms: it can either deliver the fluctuating temperature $\widetilde{u}_{\rm NN}$ in the primal form or provide the stream function $\widetilde{w}_{\rm NN}$ in the dual form. The architecture includes the periodic layer alongside one or more sequential residual layers, which utilize smooth activation functions like the hyperbolic tangent.
Figure 1: Scheme of the primal PINN architecture introduced by Jiang et al. (2023). The network inputs the microscopic coordinates $x_{1},x_{2}]$ and outputs the fluctuating temperature $\widetilde{u}_{\rm NN}$. Identical architecture is used for the dual formulation, where the output is the stream function $\widetilde{w}_{\rm NN}$.
@@ 3.2 Weak form
In this paper, we also aim to address certain limitations of PINNs for the homogenization of piecewise constant materials that are often encountered in the laminate or matrix-inclusion forms. Because the strong formulation is used in training the networks, the material function $\bm{A}(\bm{x})$ is also assumed to have continuous first-order gradients. While the material smoothness condition is not strictly necessary for the technical implementation of PINNs, ignoring it often leads to poor results (Henkes et al., 2022), so an artificial smoothing of the material function is usually employed. Moreover, in our particular homogenization setting, naively implementing piecewise-constant material coefficients leads to failure, as the network converges to a near-constant solution because the term $\nabla\cdot(\bm{A}(\bm{x}){\bm{\xi}})$ is zero in all possible collocation points. Introducing material smoothing, however, alters the underlying problem and introduces a material approximation error into the estimates of the homogenized parameter. We later show that even for smooth functions, an insufficiently gradual transition between the material phases can cause poor PINN solutions, because $\nabla\cdot(\bm{A}(\bm{x}){\bm{\xi}})$ becomes very large at the interface between the materials and very small elsewhere.
A potential remedy lies in replacing the strong form of the governing PDE with its weak form, which allows for the relaxation of the smoothness constraint in both the material function and the solution. While many potential sophisticated methods for addressing material discontinuities exist, we aim to test the limits of the standard approach by letting the VPINN have an identical architecture to the PINN described above. It is important to note that the weak formulation also broadens the options for activation functions; however, to isolate the impact of the formulation itself, we will continue to use the hyperbolic tangent activation function.
We implement the loss function introduced for Robust VPINNs introduced by Rojas et al. (2024). For the primal neural network solution $\widetilde{u}_{\rm NN}$ and a basis of test functions $\phi_{n}(\bm{x})\in H^{1}_{\rm per}(X)$, $n=1,\dots,N_{t}$ we compute the vector of residuals $\bm{r}_{\rm p}(\widetilde{u}_{\rm NN})=[r_{{\rm p}}^{1},\dots,r_{{\rm p}}^{N_{t}}]^{\rm T}$:
@@
As shown in Rojas et al. (2024), minimizing the sum of squared residuals with respect to individual test functions is equivalent to minimizing the functional
@@ 4 Numerical Experiments
This section demonstrates, on a simple example, how training PINNs in the dual formulation of the problem can complement the solution in the primal form, and compares the effectiveness of VPINNs against classic PINNs.
Consider a square domain $X$ of the size $2\pi\times 2\pi$ with a centrally placed square inclusion of the size $\pi\times\pi$ with isotropic material phases, shown in Fig. 2. For this case, the effective conductivity can be computed exactly (Obnosov, 1999) as
@@
Since the analytical solution only provides $\gamma^{\mathrm{eff}}$, we use a FEM solution for the benchmark on $\widetilde{u}$ and $\widetilde{w}$ (cf. Fig. 3). The primal and dual FEM solutions provide guaranteed upper and lower bounds to $\gamma^{\mathrm{eff}}$ (Gaynutdinova et al., 2022). We use a regular grid with triangular elements and continuous piecewise linear functions. Since the material distribution is symmetric, we only have to compute the solution for one macroscopic gradient ($\bm{\xi},\bm{\zeta}=[1,0]^{\rm{T}}$), and we will be comparing the estimates of the first diagonal element of the homogenized heat conductivity matrix only. For the solution on a regular mesh containing $128\times 128$ unique degrees of freedom (DoFs) ($129\times 129$ grid with periodic boundary nodes), these bounds are very close to the exact analytical effective conductivity (within $\pm 0.04\%$ of relative error).
In all experiments, we consider four neural network architectures of varying parameter sizes, as described in Table 1 for both primal and dual formulations. To minimise the influence of the choice of collocation points, we train all the networks on the same grid as the highest definition FEM solution (16,384 collocation points), and these points are then used for the numerical integration with the trapezoidal rule in computing the training losses of Eqs. (10)-(12) and (14), as well as effective parameter estimation in Eqs. (7) and (8).
№ neurons, periodic layer № neurons, per hidden layer № hidden layers № parameters
@@
20 20 3 1801
50 50 5 15601
Table 1: Configurations of the considered neural networks for both primal/dual and strong/weak formulations.
@@
4.1 Strong form
As discussed earlier, problems with piecewise constant material distribution cannot be solved with the classic PINNs based on the strong formulation. As a workaround, we adopt a smooth ($C^{\infty}_{\rm per}(X)$) approximation of the piecewise constant material function, similarly to Jiang et al. (2023). In this work, the approximation takes the form
$\bm{A}_{s,\varepsilon}(\bm{x})=[1-0.9\cdot p_{\varepsilon}(x_{1})\cdot p_{\varepsilon}(x_{2})]\;\bm{I},$
@@ 4.1 Strong form
$p_{\varepsilon}(x_{i})=\frac{1}{2}\left(1+\tanh\left[\frac{1}{\varepsilon}\,\sin\left(x_{i}-\frac{\pi}{2}\right)\right]\right),$
@@
To establish the link between the smoothness of the material and the solution quality, we trained all network architectures of Tab. 1 on material distributions with $\varepsilon\in\{1/10,1/20,1/30,1/40\}$ for 40,000 epochs. The networks output the primal and dual solutions, which can be used to compute estimates of the homogenized conductivity, as well as the guaranteed upper and lower bounds. For the estimates, we approximate the integrals in (7) and (8) by calculating the terms $\nabla\widetilde{u}(\bm{x})$ and $\bm{Q}\nabla\widetilde{w}(\bm{x})$ in the collocation points, and then using trapezoidal rule. Since the gradients have to be calculated during training, such primal and dual estimates provide a fast additional metric to follow during the training process. The results are summarized in Fig. 4 as dots connected by bold lines. Here we can see that these estimates can be quite precise, i.e., have a very low relative difference $(A_{1,1}^{p}-A_{1,1}^{d})/A_{1,1}^{p}$ between them (i.e., primal-dual gap), especially for the wider material transition ($\varepsilon=1/10,1/20$). However, the order of Eq. (9) is not guaranteed for these quick estimates, due to the imprecision of the integration (Haslinger and Dvořák, 1995). Furthermore, the estimates only hold for the smooth material distribution $\bm{A}_{s,\varepsilon}(\bm{x})$, which introduces its own approximation error, and the networks converge to an effective conductivity around 0.5% higher than the analytical solution.
To overcome the challenge of directly integrating the PINN solution (which is hard to integrate precisely) and to calculate the guaranteed bounds to the original problem, we can project the PINN solution in some chosen collocation points (not necessarily the training points) onto finite-element functions. Most simply, we can assume the new projected solution to have equal values to the PINN solution in the collocation points and linearly interpolate between them. Then the terms $\nabla\widetilde{u}(\bm{x})$ and $\bm{Q}\nabla\widetilde{w}(\bm{x})$ are constant on each triangular subdomain, and, for the piecewise constant material distribution and as long as the mesh is carefully constructed to align with the material phases, the primal and dual flux is constant on each triangular subdomain as well. The exact integration of such a function is then trivial.
@@
Figure 4: Comparison of the primal and dual estimates (-$\bullet$-) and guaranteed bounds (-*-) of the PINN solutions depending on the number of network parameters and the transition parameter $\varepsilon$ of the material approximation.
Generally, the primal PINN estimates converge slightly faster for an increasing number of parameters than their dual counterpart, unlike the FEM primal and dual solutions, which converge at the same rate. For smaller $\varepsilon$, larger networks are required to achieve the same primal-dual gap. The primal-dual gap decreases the fastest for the material with the most gradual transition ($\varepsilon=1/10$), at the price of wider guaranteed bounds, when the actual discontinuous material is considered. The PINN with the lowest DoFs can yield an estimate on par with the FEM benchmark if the material transition is sufficiently gradual. The best result with respect to the analytical effective conductivity is achieved for the largest network of 15,601 parameters and $\varepsilon=1/30$. On the other hand, even large networks produce poor solutions for the considered $\varepsilon=1/40$ and smaller.
The results reveal that the greatest danger in utilizing only primal PINNs for solving piecewise constant materials (or even their smooth approximations) is the inability to identify the point of failure. When solely looking at the training process (Figs. 5-6) and the resulting residual fields (Figs. 7-8), the results might not indicate poor convergence. In particular, compare training of the PINNs for $\varepsilon=1/30$ (Fig. 5) and $\varepsilon=1/40$ (Fig. 6), where the training curves look similar, but the achieved accuracy differs substantially. In the absence of the reference solution, it can also be hard to identify the point of failure, as the residuals are even lower in the area outside the material transition for the $\varepsilon=1/40$ solution than for the $\varepsilon=1/30$ solution, see Figs. 7 and 8. Only the difference between the primal and dual estimates reveals the poor quality of the PINN solutions.
@@
Figure 7: Primal solutions and the residuals of select PINNs depending on the transition parameter $\varepsilon$.
Figure 8: Dual solutions and the residuals of select PINNs depending on the transition parameter $\varepsilon$.
4.2 Weak form
The inability of the strong-form PINNs to handle discontinuous material distribution can be remedied by considering the weak form of the underlying PDE in the Variational PINNs (VPINNs). For a direct comparison of their performance to the PINNs in numerical homogenization, we consider the same network architectures as in Section 4.1, and we use the collocation points on the same $128\times 128$ grid for evaluating the integrals in Eqs. (5) and (6). As we expect the weak formulation to handle the piecewise constant material, we only train for the original, piecewise constant coefficient matrix $\bm{A}(\bm{x})$. However, in VPINNs, the choice of appropriate test functions becomes crucial (Berrone et al., 2022). Here, we investigate two types of test function bases: a spectral basis (Section 4.2.1) and a neural network basis (Section 4.2.2).
4.2.1 VPINN with spectral test functions (VSPINN)
@@
$\text{diag}\,(G^{-1}_{s,m,n})=\text{diag}\,(G^{-1}_{c,m,n})=\frac{1}{m^{2}+n^{2}}.$
@@
Here we can see that the variational PINNs with the spectral test functions (VSPINNs) can handle the discontinuity in the material and yield more precise estimates than PINNs with the small transition phase $\varepsilon=1/40$. However, larger networks are not more precise than smaller networks, and the performance of the large (1,801 and 15,601 parameters) networks with 126 test functions is even worse than the performance of the same network architectures trained with 70 test functions, especially in the primal formulation. In both cases, the training loss of the larger networks was orders of magnitude lower than the training loss of the smaller networks, see Figs. 10 and 11. This, however, does not translate into better estimates.
Because the VPINNs use the same architecture as PINNs, the pointwise strong-form residual can be calculated for comparison, shown in Figs. 12 and 13. The results suggest that the larger networks are prone to overfitting to the inprecision in the integration of the scalar products with the test functions in (5) and (6), especially considering the distortion in the solution with 15,601 parameters. This problem is worse for the basis of 126 test functions, as it contains functions with higher frequencies, for which the same integration scheme yields progressively lower precision. Integration problems might be remedied with different integration techniques, such as Filon’s quadrature (Chase and Fosdick, 1969), although these are ultimately out of the scope of this paper.
Figure 9: Comparison of the primal and dual estimates (-$\bullet$-) and guaranteed bounds (-*-) of the VSPINN solutions depending on the number of network parameters and the number of spectral test functions.
@@
Figure 12: Primal solutions and the residuals of select PINNs depending on the number of the spectral test functions $N_{t}$.
Figure 13: Dual solutions and the residuals of select PINNs depending on the number of the spectral test functions $N_{t}$.
@@
4.2.2 VPINN with neural network test functions (VNPINN)
@@
The benefits of this approach are twofold: (i) the functions are naturally from the same function space as the solution; (ii) more test functions can be added without worsening the precision of integration. The main downside is that the linear independence of these test functions is not guaranteed, mainly because of the weight-space symmetry of the neural networks (Brea et al., 2019). The Gram matrix and its inverse in (13) have to be computed numerically, and we commonly observe that for the number of test functions approaching the number of neural network parameters, this matrix is no longer positive definite, typically with a few very small negative eigenvalues. Another option would be to construct an orthogonal basis by training the networks, but this is very time-consuming in practice. Instead, when the Gram matrix was ill-condtioned, particularly for the network with 65 DoFs and $N_{t}=100$ or 200, we considered a workaround of computing the $\bm{G}$ as a diagonal matrix, where $G_{n}=\int_{X}\nabla^{\rm T}\phi_{n}\nabla\phi_{n}{\rm d}\bm{x}$.
Four differently sized PINN architectures using $N_{t}=\{50,100,200\}$ were trained. The networks with 1,801 and 15,601 parameters were additionally trained with 400 test functions. The primal and dual estimates, as well as the guaranteed bounds, are shown in Fig. 15.
@@
Here, we observe that both primal and dual solutions improve with the increasing number of parameters and test functions. Increasing the number of test functions for the smaller networks (65 and 391 parameters) has diminishing returns, as these networks perform similarly with 100 and 200 test functions. On the other hand, the precision of larger networks does not improve if the number of test functions does not increase accordingly. For example, the VNPINNs with 15,601 parameters trained with 50 test functions yield worse effective conductivity estimates and have higher strong-form residuals than the VNPINNs with 65 parameters trained with the same number of test functions. Since the achieved training loss of the large network is generally lower, the distortions in the solutions in Figs. 18 and 19 for the solution with 15,601 parameters and $N_{t}=50$ suggest the overfitting to the test functions.
@@
Figure 18: Primal solutions and the residuals of select PINNs depending on the number of the neural network test functions $N_{t}$.
@@
Figure 19: Dual solutions and the residuals of select PINNs depending on the number of the neural network test functions $N_{t}$.

### v2

@@ 3 Physically Informed Neural Networks for Homogenization
To show how existing PINN homogenization approaches can be improved by introducing the dual form training loss, we adopt the PINN architecture introduced by JIANG2023115972 (Fig. 1), which was demonstrated to produce a good approximation of the solution of (4) for a wide range of different material functions with smooth coefficients, see Fig. 1. This network is characterized by a unique periodic layer that employs learnable cosine functions. This design choice ensures that the network’s output is inherently periodic, thereby eliminating the need to incorporate a boundary condition term into the training loss.
The network takes as input the microscopic coordinates $x_{1}$ and $x_{2}$ and produces output in one of two forms: it can either deliver the fluctuating temperature $\widetilde{u}_{\rm NN}$ in the primal form or provide the stream function $\widetilde{w}_{\rm NN}$ in the dual form. Thus, by training two separate networks with the same architecture, we can obtain a primal and a dual solution to the homogenization problem. The architecture of each network includes the periodic layer alongside one or more sequential residual layers, which utilize smooth activation functions like the hyperbolic tangent.
Figure 1: Scheme of the primal PINN architecture introduced by JIANG2023115972. The network inputs the microscopic coordinates $x_{1},x_{2}]$ and outputs the fluctuating temperature $\widetilde{u}_{\rm NN}$. Identical architecture is used for the dual formulation, where the output is the stream function $\widetilde{w}_{\rm NN}$.
@@ 3.2 Weak form
In this paper, we also aim to address certain limitations of PINNs for the homogenization of piecewise constant materials that are often encountered in the laminate or matrix-inclusion forms. Because the strong formulation is used in training the networks, the material function $\bm{A}(\bm{x})$ is also assumed to have continuous first-order gradients. While the material smoothness condition is not strictly necessary for the technical implementation of PINNs, ignoring it often leads to poor results (HENKES2022114790), so an artificial smoothing of the material function is usually employed. Moreover, in our particular homogenization setting, naively implementing piecewise-constant material coefficients leads to failure, as the network converges to a near-constant solution because the term $\nabla\cdot(\bm{A}(\bm{x}){\bm{\xi}})$ is zero in all possible collocation points. Introducing material smoothing, however, alters the underlying problem and introduces a material approximation error into the estimates of the homogenized parameter. We later show that even for smooth functions, an insufficiently gradual transition between the material phases can cause poor PINN solutions, because $\nabla\cdot(\bm{A}(\bm{x}){\bm{\xi}})$ becomes very large at the interface between the materials and very small elsewhere.
A potential remedy lies in replacing the strong form of the governing PDE with its weak form, which allows for the relaxation of the smoothness constraint in both the material function and the solution. While there are many sophisticated methods for addressing material discontinuities (Liu2023; lei2025), we aim to test the limits of the standard approach by letting the VPINN have an identical architecture to the PINN described above. It is important to note that the weak formulation also broadens the options for activation functions; however, to isolate the impact of the formulation itself, we will continue to use the hyperbolic tangent activation function.
We implement the loss function for Robust VPINNs introduced by ROJAS2024116904. For the primal neural network solution $\widetilde{u}_{\rm NN}$ and a basis of test functions $\phi_{n}(\bm{x})\in H^{1}_{\rm per}(X)$, $n=1,\dots,N_{t}$ we compute the vector of residuals $\bm{r}_{\rm p}(\widetilde{u}_{\rm NN})=[r_{{\rm p}}^{1},\dots,r_{{\rm p}}^{N_{t}}]^{\rm T}$:
@@
As shown in (ROJAS2024116904), minimizing the sum of squared residuals with respect to normalized individual test functions is equivalent to minimizing the functional
@@ 4 Numerical Experiments
This section demonstrates how training PINNs in the dual formulation of the problem can complement the solution in the primal form, and compares the effectiveness of VPINNs against classic PINNs for unit cells with a singular (Section 4.1) and multiple (Section 4.2) inclusions.
4.1 Singular Inclusion
Consider a square domain $X$ of the size $2\pi\times 2\pi$ with a centrally placed square inclusion of the size $\pi\times\pi$ with isotropic material phases, shown in Fig. 2. For this case, the effective conductivity can be computed exactly (Obnosov) as
@@ 4.1 Singular Inclusion
Together with the analytical exact homogenized conductivity $\gamma^{\mathrm{eff}}$, we use a FEM solution for the benchmark on $\widetilde{u}$ and $\widetilde{w}$ (cf. Fig. 3). The primal and dual FEM solutions provide guaranteed upper and lower bounds to $\gamma^{\mathrm{eff}}$ (Gaynutdinova2022). For the sake of simplicity, we use a regular grid with triangular elements and continuous piecewise linear functions. Since the material distribution is symmetric, we only have to compute the solution for one macroscopic gradient ($\bm{\xi},\bm{\zeta}=[1,0]^{\rm{T}}$), and we will be comparing the estimates of the first diagonal element of the homogenized heat conductivity matrix only. For the solution on a regular mesh containing $128\times 128$ unique degrees of freedom (DoFs) ($129\times 129$ grid with periodic boundary nodes), these bounds are very close to the exact analytical effective conductivity (within $\pm 0.04\%$ of relative error).
In all experiments, we consider four neural network architectures of varying parameter sizes, as described in Tab. 1 for both primal and dual formulations. To reduce the impact of the choice of the collocation points, we train all the networks using the same grid as the highest resolution Finite Element Method (FEM) solution, which has 16,384 collocation points. These points are then utilized for numerical integration with the trapezoidal rule to calculate the training losses in Eqs. (19)–(21) and (23). They are also used for effective parameter estimation in Eqs. (16) and (17).
№ neurons, periodic layer № neurons, per residual layer № residual layers № parameters
@@
20 20 3 1 8011\,801
50 50 5 15 60115\,601
Table 1: Configurations of the considered neural networks for both primal/dual and strong/weak formulations, single inclusion.
@@
4.1.1 Strong form
As discussed earlier, problems with piecewise constant material distribution cannot be solved with the classic PINNs based on the strong formulation. As a workaround, we adopt a smooth ($C^{\infty}_{\rm per}(X)$) approximation of the piecewise constant material function, similarly to JIANG2023115972. In this work, the approximation takes the form
$\bm{A}_{s,\varepsilon}(\bm{x})=[1-0.9\cdot p_{\varepsilon}(x_{1})\cdot p_{\varepsilon}(x_{2})]\;\bm{I},$ (25)
@@ 4.1.1 Strong form
$p_{\varepsilon}(x_{i})=\frac{1}{2}\left(1+\tanh\left[\frac{1}{\varepsilon}\,\sin\left(x_{i}-\frac{\pi}{2}\right)\right]\right),$ (26)
@@
To establish the link between the smoothness of the material and the solution quality, we trained all network architectures of Tab. 1 on material distributions with $\varepsilon\in\{1/10,1/20,1/30,1/40\}$ for 40 00040\,000 epochs with the Adam optimizer and the starting learning rate $10^{-5}$. The networks output the primal and dual solutions, which can be used to compute estimates of the homogenized conductivity, as well as the guaranteed upper and lower bounds. For the estimates, we approximate the integrals in (16) and (17) by calculating the terms $\nabla\widetilde{u}(\bm{x})$ and $\bm{Q}\nabla\widetilde{w}(\bm{x})$ in the collocation points, and then using trapezoidal rule. Since these gradients must be calculated during training, the gap between the primal and dual estimates provides a fast additional convergence metric to follow during the training process. The final estimates are summarized in Fig. 4 as dots connected by bold dashed lines. Here we can see that these estimates can be quite precise, i.e., have a very low relative difference $(A_{1,1}^{p}-A_{1,1}^{d})/A_{1,1}^{p}$ between them (i.e., primal-dual gap), especially for the wider material transition ($\varepsilon=1/10,1/20$). However, the order of Eq. (18) is not guaranteed for these quick estimates, due to the imprecision of the integration (haslinger1995optimum). Furthermore, the estimates only hold for the smooth material distribution $\bm{A}_{s,\varepsilon}(\bm{x})$, which introduces its own approximation error, and the networks converge to an effective conductivity around 0.5% higher than the analytical solution.
In principle, the computation of the guaranteed bounds requires two conditions to be satisfied: (i) the integrals in Eqs. (16) and  (17) are calculated precisely; and (ii) the approximate solutions $\widetilde{u}(\bm{x})$ and $\widetilde{w}(\bm{x})$ are from the subspaces of $H^{1}_{\rm per}(X)$. To tackle the challenge of precisely integrating the PINN solution (since it can be difficult to derive a closed form), we project it onto finite-element functions, which also form a subspace of $H^{1}_{\rm per}(X)$. The simplest approach is to set the projected solution in the nodes of some finite element mesh to be equal to the PINN values at these points and interpolate linearly between them. As a result, the terms $\nabla\widetilde{u}(\bm{x})$ and $\bm{Q}\nabla\widetilde{w}(\bm{x})$ remain constant within each triangular subdomain. If the material distribution is piecewise constant and the mesh is designed to align with the material phases, both the primal and dual flux will also be constant in each triangular subdomain. This makes the exact integration of the function trivial. Here, we use the same regular mesh that was used for the FEM benchmark computations, which also aligns with the training collocation points.
@@
Figure 4: Comparison of the primal and dual estimates (-$\bullet$-) and guaranteed bounds (-$\star$-) of the PINN solutions depending on the number of network parameters and the transition parameter $\varepsilon$ of the material approximation.
Generally, the primal PINN estimates converge slightly faster for an increasing number of parameters than their dual counterpart, unlike the FEM primal and dual solutions, which converge at the same rate. For smaller $\varepsilon$, larger networks are required to achieve the same primal-dual gap. The primal-dual gap decreases the fastest for the material with the most gradual transition ($\varepsilon=1/10$), at the price of wider guaranteed bounds, when the actual discontinuous material is considered. The PINN with the lowest number of DoFs can yield an estimate on par with the FEM benchmark if the material transition is sufficiently gradual. The best result with respect to the analytical effective conductivity is achieved for the largest network of 15,601 parameters and $\varepsilon=1/30$. On the other hand, even large networks produce poor solutions for the considered $\varepsilon=1/40$ and smaller.
The results reveal that the greatest danger in utilizing only primal PINNs for solving piecewise constant materials (or even their smooth approximations) is the inability to identify when the achieved solution is poor. When solely looking at the training process (Figs. 56) and the resulting residual fields (Figs. 78), the results might not indicate poor convergence. In particular, compare training of the PINNs for $\varepsilon=1/30$ (Fig. 5) and $\varepsilon=1/40$ (Fig. 6), where the training curves look similar, but the achieved parameter estimates differ substantially. In the absence of the reference solution, it can also be hard to identify the point of failure, as the residuals are even lower in the area outside the material transition for the $\varepsilon=1/40$ solution than for the $\varepsilon=1/30$ solution, see Figs. 7 and 8. Only the difference between the primal and dual estimates reveals the poor quality of the PINN solutions.
@@
Figure 7: Primal solutions (top) and the residuals (bottom) of select PINNs depending on the transition parameter $\varepsilon$.
Figure 8: Dual solutions (top) and the residuals (bottom) of select PINNs depending on the transition parameter $\varepsilon$.
4.1.2 Weak form
The inability of the strong-form PINNs to handle discontinuous material distribution can be remedied by considering the weak form of the underlying PDE in the Variational PINNs (VPINNs). For a direct comparison of their performance to the PINNs in numerical homogenization, we consider the same network architectures as in Section 4.1.1, and we use the collocation points on the same $128\times 128$ grid for evaluating the integrals in Eqs. (14) and (15). As we expect the weak formulation to handle the piecewise constant material, we only train for the original, piecewise constant coefficient matrix $\bm{A}(\bm{x})$. However, in VPINNs, the choice of appropriate test functions becomes crucial (Berrone2022). Here, we investigate two types of test function bases: a spectral basis (paragraph VSPINN below) and a neural network basis (paragraph VNPINN below).
VPINN with spectral test functions (VSPINN)
@@ 4.1.2 Weak form
$\text{diag}\,(G^{-1}_{s,m,n})=\text{diag}\,(G^{-1}_{c,m,n})=\frac{1}{m^{2}+n^{2}}.$ (27)
@@
Here we can see that the variational PINNs with the spectral test functions (VSPINNs) can handle the discontinuity in the material and yield more precise estimates than PINNs with the small transition phase $\varepsilon=1/40$. However, larger networks are not more precise than smaller networks, and the performance of the large (1 8011\,801 and 15 60115\,601 parameters) networks with 126 test functions is even worse than the performance of the same network architectures trained with 70 test functions, especially in the primal formulation. In both cases, the training loss of the larger networks was orders of magnitude lower than the training loss of the smaller networks, see Figs. 10 and 11. This, however, does not translate into better estimates.
Because the VPINNs use the same architecture as PINNs, the pointwise strong-form residual can be calculated for comparison, shown in Figs. 12 and 13. The results suggest that the larger networks are prone to overfitting to the inprecision in the integration of the scalar products with the test functions in (14) and (15), especially considering the distortion in the solution with 15 60115\,601 parameters. This problem is worse for the basis of 126 test functions, as it contains functions with higher frequencies, for which the same integration scheme yields progressively lower precision. Integration problems might be remedied with different integration techniques, such as Filon’s quadrature (Chase), although these are ultimately out of the scope of this paper.
Figure 9: Comparison of the primal and dual estimates (-$\bullet$-) and guaranteed bounds (-$\star$-) of the VSPINN solutions depending on the number of network parameters and the number of spectral test functions.
@@
Figure 12: Primal solutions and the residuals of select VSPINNs depending on the number of the spectral test functions $N_{t}$.
Figure 13: Dual solutions and the residuals of select VSPINNs depending on the number of the spectral test functions $N_{t}$.
@@
VPINN with neural network test functions (VNPINN)
@@
The benefits of this approach are twofold: (i) the functions are naturally from the same function space as the solution; (ii) more test functions can be added without worsening the precision of integration. The main downside is that the linear independence of these test functions is not guaranteed, mainly because of the weight-space symmetry of the neural networks (brea2019). The generally non-diagonal Gram matrix and its inverse in (22) have to be computed numerically, and we commonly observe that for the number of test functions approaching the number of neural network parameters, this matrix is no longer positive definite, typically with a few very small negative eigenvalues. Another option would be to construct an orthogonal basis by training the networks, but this is very time-consuming in practice. Instead, when the full Gram matrix was ill-conditioned, particularly for the network with 65 DoFs and $N_{t}=100$ or 200, we set all its non-diagonal values to zero.
Four differently sized PINN architectures using $N_{t}=\{50,100,200\}$ were trained. The networks with 1 8011\,801 and 15 60115\,601 parameters were additionally trained with 400 test functions. The primal and dual estimates, as well as the guaranteed bounds, are shown in Fig. 15.
@@
Here, we observe that both primal and dual solutions improve with the increasing number of parameters and test functions. Increasing the number of test functions for the smaller networks (65 and 391 parameters) has diminishing returns, as these networks perform similarly with 100 and 200 test functions. On the other hand, the precision of larger networks does not improve if the number of test functions does not increase accordingly. For example, the VNPINNs with 15 60115\,601 parameters trained with 50 test functions yield worse effective conductivity estimates and have higher strong-form residuals than the VNPINNs with 65 parameters trained with the same number of test functions. Since the achieved training loss of the large network is generally lower, the distortions in the solutions in Figs. 18 and 19 for the solution with 15 60115\,601 parameters and $N_{t}=50$ suggest the overfitting to the test functions. While this negative effect is less pronounced when the number of test functions is large (i.e., $N_{t}\in{200,400}$), the comparison between the training loss and the primal-dual gap in Figs. 16 and 17) shows that while the convergence in training loss has not been reached in the allocated 40 00040\,000 epochs, the primal-dual gap has reached convergence around 30 00030\,000-th epoch, which, again, suggests the overfitting to the test functions.
@@
Figure 18: Primal solutions (top) and the residuals (bottom) of select VNPINNs depending on the number of the neural network test functions $N_{t}$.
@@
Figure 19: Dual solutions (top) and the residuals (bottom) of select VNPINNs depending on the number of the neural network test functions $N_{t}$.